Status Report

NASA Space Station On-Orbit Status 9 Jan 2004

By SpaceRef Editor
January 9, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 9 Jan 2004

All ISS systems continue to function nominally, except as noted previously or below.

Update on cabin pressure drop (ref. Status report 1/7):

Cabin pressure decay this morning was at a low 0.8 lb/day (nominal station leak rate: ~0.08 lb/day), with total pressure at 14.0 psi (723.8 mmHg [or “torr”, after Torricelli]).  Since some hardware is not certified at 13.9 psi or lower, ISS will be repressurized on reaching 13.91 psi (early Monday morning, 1/12), probably using U.S. nitrogen (N2) from the high-pressure tanks (HPTs) outside the Airlock to go back up to 14.3 psi (requiring 25 lb N2, from 307 lb available).  C&W (caution & warning) resets may also be required.  Crew is continuing checkout of on-board systems, which includes “nonintrusive” module isolations (on Airlock (A/L), DC-1, Soyuz and Progress).  By end of this weekend, all most-likely components and modules will have been checked out.  Plans for isolating the Russian (RS) from the U.S. (USOS) segment are being developed jointly, to be implemented next Wednesday, 1/14 (if still required then).  This will necessitate moving some equipment to the RS, since the crew would reside there for the duration of the isolation (4-5 days).  Since the OpsLAN (operations local area network) cable between segments would have to be disconnected, some computer/data capabilities would not be available.  Workarounds are being set up where really necessary.

FE Alexander Kaleri worked with a ground specialist on the Vozdukh carbon dioxide (CO2) scrubber, as part of the on-going search for the cause of the minute cabin pressure drop (see above).  Other Russian systems with vacuum-connections to be checked for this purpose over the weekend are the BMP (harmful impurities filtration system), two pneumatic panels (PK2, PK3) of the TCS (thermal control system) recharge loop, and the SKV-2 air conditioner, which had been evacuated during the repair activities on 12/16-17.  [Currently done: Vozdukh and BMP, both negative.]

Kaleri deactivated the IK0501 gas analyzer (GA) in the Service Module (SM) and exchanged its BF carbon dioxide (CO2) filter assembly with a new unit from the FGB.  GA was reactivated and the spent BF stowed for disposal (replaced last: 11/29). [IK0501 is an automated system for measuring CO2, O2, and H2O in the air as well as the flow rate of the gas being analyzed.]

Sasha also took the periodic reading of the cabin air’s current CO2 partial pressure in the SM using the U.S. CDMK (CO2 monitor kit), for calldown to MCC-Houston (along with the battery status) for use in trending analyses.  The CDMK was then restowed at the SM Central Post.

The Elektron oxygen generator is currently up and running (16A mode), after Kaleri performed the scheduled troubleshooting activities, opening the BZh Liquid Unity of the GZhS air/liquid separator system to adjust a pressure setting device, then restarting the unit.

The crew burned two more SFOG (solid-fuel oxygen generator) candles today, one at 6:35am EST, the other at 9:55am.  This brings the number of candles that have so far been used since 12/31 to 18 (2 units/day).  [Of the remaining 124 SFOGs (Russian: TGK), three have failed to ignite in their generators, and a fourth has a damaged seal and is unusable.  The onboard TGK supply comes from two batches, both of which have reached the end of their original certified life, but all evaluations and reviews necessary for continued use of the TGKs are complete.]

Foale continued payload activities on the PFMI (Pore Formation & Mobility Investigation) in the MSG (Microgravity Science Glovebox).  [After removing and stowing sample PFMI-15 from the thermal chamber, Mike exchanged the PFMI Microdrive Assembly and removed the video tape made yesterday.  He was thanked for yesterday’s successful run.]

Alex performed the regular once-a-week maintenance reboot on the operational PCS laptops and also restarted the OCA comm router laptop (every two weeks).

The crew worked out on TVIS, RED exerciser and VELO cycle with load trainer.

The CDR completed the regular task of transferring data files from the physical exercise equipment to the MEC (medical equipment computer) via memory card and RED log entries, for downlink on OCA comm.  Later, he performed the periodic transfer of accumulated data files from the wristband HRM (heart rate monitor) receiver stations to the MEC for downlink, then deleting them on the HRM.  

Mike also conducted the regular routine status checkup on autonomously running Increment 8 payloads in the Lab, while Sasha Kaleri prepared the daily IMS delta file for automatic export/import to update the database.

The FE also conducted his regular weekly IMS (inventory management system) tagup with IMS specialists at MCC-M, discussing open issues concerning identification of equipment and storage locations for the IMS databases.  [Today’s issues involved location and status of SUD components removed from the FGB on 12/29 and 12/30, the estimated available space behind FGB panel 312 for cargo to be delivered by Progress 13P, an upcoming (1/15) audit of RS stowage locations, and others.]

The crew started their second session of the NASA/JSC renal (kidney) stone experiment activities.  Mike unstowed and set up the experiment hardware and performed an inspection of the available UCDs (urine collection devices) for damage and an inventory of the remaining equipment.  Both he and Sasha Kaleri began the first day of this session by starting their diet logs, which record all foods and drinks consumed during the next 48 hours.   [This long-range preventive medicine investigation features daily random ingestion of either potassium citrate or placebo tablets.  It is Dr. Peggy Whitson’s double-blind research study investigating methods to prevent formation of kidney stones in zero-G.  Part of the experiment consists in keeping a metabolic diet log (food and fluid intake), followed by collection of urine samples several times per day during each week-long session.]

At 10:45am, the CDR had his weekly teleconference with the JSC Astronaut Office (Kent Rominger), followed by the crew’s weekly teleconference with the ISS Flight Director at MCC-H.

The scheduled retransfer of A/L cargo items from their temporary stowage in the Node to the A/L was deferred.  [The equipment had to be cleared from the A/L on 1/5 for Mike’s EMU swap/resizing activities.]

Yesterday’s station reboost by Progress 12 P was successfully completed as scheduled, with parameters being as expected.  Approximately 72 kg of propellants were expended, for a delta-V of 1.4 m/s.  Mean orbit altitude was raised by 2.5 km (1.6 miles).

After the reboost, the ISS maneuvered from the earth-oriented LVLH (local vertical/local horizontal) attitude to sun-pointing XPOP (x-axis perpendicular to orbit plane).  The solar array BGAs (beta gimbal assemblies) 2B and 4B were moded to Autotrack, biased 45 degrees for drag reduction.  [While flying in XPOP, the Lab’s nadir window, used for science, will be within 85 deg of the velocity vector (flight direction) for about half the orbit.  To protect it against particle impacts, its shutter must remain closed during this period, i.e., from 1 minute after orbital noon to 1 minute before orbital midnight.]

Today’s CEO targets, in the current XPOP attitude constrained by flight rule to fewer near-vertical targets due to shutter closure, wereSaharan dust (Dynamic event.  Blowing dust is reducing visibilities at stations near Lake Chad),Africa’s Great Lakes(Dynamic event. Fair weather as ISS passed along the west side of Lake Tanganyika and then across Lake Malawi),Bamako, Mali (nadir pass over this hard-to-see capital city:  looking for a devegetated circular area on the Niger River with radiating transportation spokes), andPatagonian Glaciers(two overflight opportunities.  With good focus, 400- and 800-mm lens images have started to reveal excellent detail of moraine structure and crevasse patterns on glacier surfaces.  Most of the remote Patagonian glaciers have been imaged at shorter focal lengths so that detailed (longer-lens) images were now requested.  Weather is clearing from the west, to reveal the usually obscured Pacific coast.  The glaciers are distinct features in the narrow zone between the coastline and the spine of the Andes Mountains).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 2:35pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is poweredOn.  Vozdukh CO2 scrubber is On (in Manual Mode 3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off (repair incomplete).
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 26.6; ppO2 (mmHg) — 148.7; ppCO2 (mmHg) — 3.8;
  • SM Transfer Compartment:  Pressure (mmHg/psi) — 742; temperature (deg C) — 19.5.
  • FGB Cabin:  Pressure (mmHg/psi) — 744; temperature (deg C) — 22.7.
  • Node:  Pressure (mmHg/psi) — 727.8; temperature (deg C) — 24.1 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg/psi) — 726.84; temperature (deg C) — 24.7; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg/psi) — 727.04; temperature (deg C) — 23.9; shell heater temp (deg C) — 24.2, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 24.7
  • PMA-2:  Shell heater temp (deg C) — 13.9.

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Autotrack (suntracking, biased 45 deg. for drag reduction). 
  • SM batteries:  Battery #8, formerly known as #7, is still disconnected in slot #8 for troubleshooting; battery #4 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.  
  • FGB batteries:  Battery #4 is off (capacity restoration mode, ROM); battery #6 is in “Cycle” mode; all other batteries (4) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available:3637 kg (8018 lb) as of 1/1/04 [SM(755) + FGB(2530) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH  XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9.4 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at MBS PDGF #1/LEE B, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 6:34am EST [= epoch]):

  • Mean altitude — 371.4 km
  • Apogee — 376.9 km
  • Perigee — 365.8 km
  • Period — 92.0 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0008258
  • Orbits per 24-hr. day — 15.66
  • Mean altitude gain last 24 hours — 2.5 km
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 29326

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.